MARS DESERT RESEARCH STATION

Crew Photos – April 18th

Convoy.

 

A bit of geology.

 

Astronaut.

 

Com experiment.

 

Looking at Mars.

 

Hello from Mars.

EVA Report – April 18th

Weather: sunny, and very windy
Temperature: around 20°C
Location: UTM NAD27 4253000 – 519000
Duration: 120 min
Team: Aurian d’Avernas (Crew Commander), Calogero Montedoro (Crew Biologist), Mathieu Vander Donckt (Crew Journalist)
Purpose:
–       Collecting new soil samples
–       Telecommunication: installation of omni-and-unidirectional telecommunication relays
Outputs:
–       Collecting succeeded.
–       Telecommunication relays installation failed due to too heavy winds. Reported to tomorrow.

Commander Report – April 18th

Today was a rather common day in the schedules. Breakfast at 8 am, followed by an EVA from 9 to 11:59 am with Mathieu, Calogero and myself to install my telecommunication relays and continue signal tests and recordings, and to collect new soil samples for the Mathieu’s experiment, together with the logistic help of Calogero. Unfortunately, winds were too strong to work properly on the top of the hills with the telecommunication relays. This part of EVA is then reported to tomorrow. About Mathieu, he was still able to collect his new soil samples. Back from EVA, we enjoyed culinary talents of Calogero to taste rice with vegetables and a home-made pizza. During the afternoon, every crew member continued to work on his experiment in the Hab for the last EVAs of the next days.

Sol Summary – April 18th

MDRS Daily Summary Report for sol 8

Summary Title: 9th EVA

Mission Status:  Good

Sol Activity Summary:

Hab
8.00:   Breakfast
10.00: Muscle wasting exercises
13.00: Lunch

GreenHab
/

ScienceDome
Experiment reports redaction

EVA
10.30: UTM NAD27 12S 4251500-518500
Telecommunication relays test
Diffractometer study: data collection

Weather:
Sunny and strong wind

Crew Physical Status
Everyone is fine

Look Ahead Plan:
EVA + Experiments

Reports to be file: Operation Report, Daily Summary Report, Commander Report, EVA 9 Report

Support Requested: